OpenStudy (anonymous):

If y varies directly as x, and y = 9 when x=2, what is y when x = -3? I understand that the equation is step up this way, y = kx and the number are plugged in, but my answer is not the same as the answer key. Thx

OpenStudy (saifoo.khan):

What value u got for k?

OpenStudy (anonymous):

9/2 = K

OpenStudy (saifoo.khan):

\[\Huge \checkmark\]

OpenStudy (saifoo.khan):

\[\Huge y = \frac{\frac92}{-3}\]

OpenStudy (saifoo.khan):

\[y = \frac92 \div -3\] \[y = \frac92 \times -\frac13\] Solve now.

OpenStudy (anonymous):

Maybe my answer key has the wrong answer of 29

OpenStudy (saifoo.khan):

Yes it does.

OpenStudy (anonymous):

thanks

OpenStudy (saifoo.khan):

no problem.